a rectangular rug has an area of 96 square feet and a perimeter of 4o feet. The length is grater than the width. Suppose the width is x feet. Use the fact that the perimeter is 40 feet to find an expression for the light as a function of x. Find the length and width of the rug.
Answer by Boreal(15235)

width is x
length is 20-x, so half perimeter is 20 (L+W).
Area is x(20-x) and that equals 96
therefore 20x-x^2=96
x^2-20x+96=0
(x-12)(x-8)=0
x=12 feet or 8 feet.
Use 8 feet for width since given that width is smaller.
It is 8 ft x 12 ft.
That is an area of 96 ft^2.